crystaldavis12:

A school spent $150 on advertising for a breakfast fundraiser. Each plate of food was sold for $8.00 but cost the school $2.00 to prepare. After all expenses were paid, the school raised $2,400 at the fundraiser. Which equation can be used to find x, the number of plates that were sold?

Mercury:

choices: 6x – 150 = 2,400 6x + 150 = 2,400 10x – 150 = 2,400 10x + 150 = 2,400

Mercury:

"after all expenses were paid, the school raised 2,400" so, (revenue) - (expenses) = 2,400 revenue is the (profit per plate)(number of plates sold). fill in the appropriate coefficient and variable from the problem expenses = advertising revenue = $150 once you have all that, plug revenue and expenses back into the equation, then see which answer choice matches